# Survey on Hand Gesture Recognition from Visual Input

### v1

@@ I Introduction
Compared to other recent surveys in the field, this survey offers a more comprehensive yet focused review of research in the field of Hand Gesture Recognition (HGR) using visual input. Unlike [150] that explores various modalities beyond visual input (e.g., EEG, EMG, and audio) with a primary emphasis on Sign Language Recognition (SLR) and limited discussion on hand gesture estimation, our work specifically targets the visual input-based approaches. We categorize HGR into distinct tasks such as classification and estimation, as well as methodological distinctions like single-camera versus multi-camera setups, offering deeper insights into both general HGR and specialized topics like body reconstruction with an emphasis on HGR. Compared to [118] that categorizes Human Motion Recognition (HMR) methods broadly into vision sensor-based (VS) and wearable sensor-based (WS) ones, and further divide HGR into markerless and marker-based approaches, our survey focuses exclusively on visual input-based methods, ensuring a more detailed exploration of modern challenges, datasets, and trends. Moreover, we include research from 2018 to present, incorporating all latest advancements, challenges and trends. This focus and up-to-date analysis allow our survey to present a more relevant and actionable synthesis of current and future directions in HGR research.
@@ II-C Multimodal Gestures
II-D Hand Gesture Recognition
@@ III-A Article Selection
The selection of articles for this study targets all the recent papers in the computer vision field that specifically focus on hand gesture recognition from images and videos. To ensure a comprehensive review, we began by identifying the top-ranked venues in computer vision as listed under the ”Top Publications - Computer Vision & Pattern Recognition” section on Google Scholar [^1] https://scholar.google.com/citations?view_op=top_venues&hl=en&vq=eng_computervisionpatternrecognition. Using a crawler developed for this purpose222the code will be made publicly available when the article will be published, we retrieved 2,307 articles from these top venues. Subsequently, we conducted a systematic screening process in which the titles and abstracts of all articles were carefully reviewed based on predefined inclusion criteria detailed in Section III-B, which prioritized recent studies focusing on hand gesture recognition techniques, systems, or applications. This rigorous filtering ensured that only the most relevant to hand gesture recognition contributions from Google Scholar’s top venues for computer vision were included in the survey, and resulted in 37 articles that were directly relevant to the task. Given the limited number of relevant studies (from Google Scholar), we widened the search base by conducting additional searches in all venues listed in Scopus using specific boolean search queries, which are listed in Table I. The queries were carefully designed to contain the most relevant keywords to hand gesture recognition. To ensure consistency with the Google Scholar approach, we systematically reviewed the titles and abstracts of the matching papers from each query, assessing their alignment with the inclusion/exclusion criteria outlined in Section III-B. Articles were excluded if they lacked relevance to the hand gesture recognition task, its techniques, systems, or applications.
@@
”RGB” AND (”hand” OR ”gesture”) AND (”recognition” OR ”classification”) 1476 910 19
”sign” AND ”gesture” AND ”recognition” 2668 1442 29
”Hand” AND (”gesture” OR ”pose”) AND (”recognition” OR ”estimation”) AND (”RGB” OR ”video” OR ”skeleton” OR ”multi modal”) 2428 746 27
”Multiview” AND ”Camera” AND ”Hand” AND ”Gesture” 5 1 1
TABLE I: Overview of queries addressed to Scopus database and paper counts
Table I includes the following columns:
@@
To determine the values in these columns, we employed a systematic methodology to curate the Scopus papers. The queries were executed sequentially, starting with the first query at the top of the table and progressing to the last query at the bottom. The process began with a broad query focused on terms such as ”RGB,” ”hand,” ”gesture” and related recognition tasks, providing a foundation for further exploration. This was followed by an expansion into more specific domains, such as sign language recognition, addressed by the second query. To avoid redundancy, articles retrieved from the first query were excluded from the results of the second query. Similarly, the third query broadened the search to include tasks related to hand pose estimation and multimodal approaches, while ensuring duplicates from previous queries were removed. For each collection of retrieved papers, topic modeling using Non-Negative Matrix Factorization (NNMF) was performed in a similar way as detailed in Section III-D. Specifically, NNMF was applied to the titles, keywords, and abstracts of the papers to extract thematic topics. Articles with topics related to the research focus were retained, significantly narrowing down the number of candidate papers. Finally, a selection process was performed based on the inclusion/exclusion criteria outlined in Section III-B, resulting in the final set of papers used in this study.
@@ III-B Inclusion/Exclusion Criteria
  • Studies were included only if published between 2018 and 2024,
@@ III-C Quantitative Analysis
The search and filtering methodology described above helped us identify 113 studies (37 from Scholar and 76 from Scopus) that propose efficient methodologies for hand gesture recognition. To better understand how these studies are distributed over time and across different forums we performed a quantitative analysis, which is summarized in Table II and Fig. 3. The visualizations provide a comprehensive overview of the research landscape of hand gesture recognition, revealing trends in publication frequency, preferred venues, and topics.
@@ III-C1 Distribution of Articles to Venues
As shown in Table II, the main forums where research in hand gesture recognition has been published include the major computer vision conferences such as CVPR, ICCV, ICCVW, ECCV, and the International Journal of Computer Vision. These forums list more than 50% of the relevant works that contribute to advancements in this field.
@@ III-C2 Research Timeline by Topic
As shown in Fig. 3 the main topics discussed in the articles. The methodology for extracting the topics from the collection of articles is detailed in Section II-D that follows. The bar charts depict the annual occurrences of individual topics, while the line plot represents the total number of publications per year. The analysis reveals a substantial increase in publications in 2024, signaling growing research interest and advancements in the field of hand gesture recognition.
@@ IV-C Classify by Hand Gesture capture and representation
Figure 8: Hand Capture Methods: Skeleton-based (left image) and box/filter-based (right image).
@@ IV-C1 Skeleton-Based Captures
Skeleton-based methods represent the hand using a skeletal structure that consists of key joints and connections [149, 33, 105]. These captures typically utilize graphs, where nodes represent specific hand joints or segments such as fingers, and edges denote the relationships or connections between them. This graph-based representation enables the system to focus on the underlying structure and movement dynamics of the hand. By abstracting the hand into a simplified skeleton, these methods are particularly effective in scenarios where the precise articulation of fingers and joints is important.
@@ IV-C2 Box/Filter-Based Captures
Box/filter-based methods follow a different approach, focusing on capturing the hand’s region using bounding boxes or regions of interest [4, 24, 131]. These methods involve detecting the hand from the image using a bounding box and once the hand is isolated, applying additional filters or algorithms to extract meaningful features related to the gesture, such as shape, texture, or movement patterns.
Skeleton-based captures excel in scenarios that require precise motion analysis and are robust against background noise. However, these methods often require more computational resources to represent the hand as a graph, making them more suitable for applications where accuracy and detailed articulation are prioritized over speed or simplicity. In contrast, box/filter-based methods are well-suited for working with 2D data and are easily adaptable to a wide range of environments. These methods are generally more user-friendly and are widely applied in hand gesture classification tasks (as described in Section IV-F, where Bayes’ Theorem demonstrates that box/filter methods are predominantly used in classification tasks), where the primary goal is to categorize gestures, rather than in estimation tasks, which require detailed hand representation. Despite their advantages, box/filter-based methods are sensitive to external factors such as lighting conditions, background clutter, and variations in hand appearance, which can negatively impact recognition performance.

### v3

@@ I Introduction
Figure 2: Timeline of the main advances in HGR for the last 40 years.
The field of hand gesture recognition (HGR) has undergone a significant evolution over the years, transitioning from early reliance on sensor-based data gloves, exemplified by the Sayre Glove in 1977 and the Digital Entry Data Glove in 1983, to increasingly sophisticated vision-based approaches, with the first such systems appearing in the early 1980s and a "true" vision-based approach reported by in 1993 [161]. Key advancements included the development and application of various feature extraction techniques such as image moments, convex hull analysis, Gabor filters, Histogram of Oriented Gradients [32], and Local Binary Patterns [147]. Furthermore, the application of classical machine learning algorithms like Hidden Markov Models in the late 1990s, k-Nearest Neighbors with foundational early applications in HGR in the early 2000s, Naive Bayes and Support Vector Machines applied to HGR in the early 2000s, played a crucial role in enabling gesture classification. This rich history of foundational work provided the essential knowledge and techniques that paved the way for the more recent and rapid progress observed in the last five years. In this later period the transformative impact of deep learning methodologies, fueled by increased computational power and the availability of larger datasets, led to significant breakthroughs in the accuracy and robustness of hand gesture recognition systems during that time. Fig. 2 provides a timeline of the evolution of HGR over the years.
@@
Existing surveys in the field, their scope, methodology, contribution, and novelties are summarized in Table I.
TABLE I: An overview of recent surveys in the field and their features.
Article Scope Methodology Contribution Novelty Limitations
A systematic review on hand gesture recognition techniques, challenges and applications (2019) [225] Research on HGR. Surveys 244 papers (49 excluded). Compares techniques, applications, and challenges. Discusses gesture acquisition methods (image-based, glove-based, motion sensing). Focuses on the latest research comparisons across multiple dimensions. Limited by its 2016-2018 timeframe, reliance solely on the IEEE Xplore database. Lacks current and visual-input-focused scope.
Hand Gesture Recognition Based on Computer Vision (2020) [149] Computer vision techniques for HGR. Reviews literature on HG techniques. Tabulates performance, classification, datasets, and camera types. Lists merits/limitations of methods. Comparative review of studies using computer vision. First review to analyze camera types, distance limitations, and recognition rates. Published in 2020, it lacks latest advancements, a systematic methodology for paper selection, and focus on hand pose estimation.
Methods, databases and recent advancement of vision-based HGR for HCI systems (2021) [176] Vision-based methods for human-computer interaction, focusing on static/dynamic gestures and dataset characteristics. Critical analysis of 85+ papers (2016-2021). Compares preprocessing techniques, feature descriptors, and classifiers across 10 benchmark datasets. Detailed comparison of recognition accuracy across illumination conditions. Introduces HCI-specific evaluation metrics beyond pure accuracy. First survey to explicitly link gesture recognition performance to HCI usability requirements and ergonomic factors. Primarily HCI-focused and published in 2021, it misses recent HGR advancements, datasets, and the latest systematic method classification.
A Structured and Methodological Review on Vision-Based Hand Gesture Recognition System (2022) [5] Vision-driven hand gesture recognition across various camera orientations. Literature survey of 108 articles summarizes and compares similar methodologies. Identifies limitations in image acquisition, segmentation, tracking, feature extraction, and classification. Aims to identify improving areas and underdeveloped challenges. Covers up to 2022, missing recent works and the latest systematic dataset/methodology synthesis.
A Systematic Review of Hand Gesture Recognition (2024) [63] Comprehensive analysis of HGR advancements from 2018-2024, including deep learning and sensor fusion approaches. Systematic review using PRISMA framework. Analyzes 127 papers. Classifies by sensing modality and application domain. Identifies emerging trends and persistent challenges (cross-cultural gestures, real-time deployment). First review to systematically benchmark post-2020 deep learning architectures and their deployment constraints in HGR. Includes sensor-based methods, not solely visual input; less comprehensive on specific visual techniques classification (e.g., box-filter vs. skeleton).
A Decade of Progress in Human Motion Recognition (2024) [142] HMR covering vision sensor-based (VSM) and wearable sensor-based methods (WSM). Categorization of research into VSM and WSM, assessed based on sensors, classification algorithms, datasets, gesture types, target body parts, and performance. Provides a comprehensive and systematic understanding of the latest developments in HMR techniques from 2010 to 2020. Categorizes HMR Assesses HMR from multiple viewpoints to provide a comprehensive overview of research trends and technological advancements. Broader HMR scope (not just HGR), 2010-2020 timeframe, and inclusion of wearable sensors limits its HGR visual-input focus.
A Methodological and Structural Review of HGR (2024) [183] HGR techniques and data modalities. Reviews HGR techniques and modalities (2014–2024). Assesses efficacy through recognition accuracy. Comprehensive review of sensor tech. Identifies gap in continuous gesture recognition. First to comprehensively review HGR data modalities within this timeframe. Covers diverse data modalities, not exclusively visual input, potentially weakening focus on visual-input HGR.
Survey on vision-based dynamic hand gesture recognition (2024) [200] Vision-based dynamic HGR. Summarizes techniques, merits/demerits, datasets, accuracy, and algorithms. Scrutinizes traditional vs. deep learning. Examines performance of traditional and deep learning methods. First to systematically compare traditional and deep learning for dynamic gestures. Focuses on dynamic HGR; less dedicated coverage of static gestures or 3D hand pose estimation.
Survey on Hand Gesture Recognition from Visual Input. (2025) [Current survey] Recent advancements in hand gesture and 3D hand pose recognition from RGB, depth, and video data. Systematic review of 137 papers from top computer vision venues using topic modeling. Classifies methods by task, input, capture, and ML techniques. Comprehensive overview of recent research, solutions, and datasets. Organized presentation of research approaches. Review of current research trends. Addresses a gap in comprehensive coverage of recent developments. More focused review concentrating on visual input.
@@
Compared to other recent surveys in the field, this survey offers a more comprehensive yet focused review of research in the field of Hand Gesture Recognition (HGR) using visual input. Unlike [183] that explores various modalities beyond visual input (e.g., EEG, EMG, and audio) with a primary emphasis on Sign Language Recognition (SLR) and limited discussion on hand gesture estimation, our work specifically targets the visual input-based approaches. We categorize HGR into distinct tasks such as classification and estimation, as well as methodological distinctions like single-camera versus multi-camera setups, offering deeper insights into both general HGR and specialized topics like body reconstruction with an emphasis on HGR. Compared to [142] that categorizes Human Motion Recognition (HMR) methods broadly into vision sensor-based (VS) and wearable sensor-based (WS) ones, and further divide HGR into markerless and marker-based approaches, our survey focuses exclusively on visual input-based methods, ensuring a more detailed exploration of modern challenges, datasets, and trends. Moreover, we include research from 2018 to the present, incorporating all the latest advancements, challenges, and trends. This focus and up-to-date analysis allow our survey to present a more relevant and actionable synthesis of current and future directions in HGR research.
@@ II-C Multimodal Gestures
II-D Hand Gesture Recognition (HGR)
@@
II-E Unsupervised and Self-Supervised Learning in HGR
Beyond traditional supervised approaches that rely on extensively labeled datasets, unsupervised and self-supervised learning paradigms are used in HGR, especially for scenarios with limited annotated data [24]. These methods aim to learn meaningful representations or to discover patterns from visual input without manual labeling, thereby reducing annotation costs and enabling models to leverage vast amounts of readily available unlabeled hand gesture data.
Unsupervised learning in HGR focuses on identifying inherent structures within unlabeled gesture data. This can manifest as clustering visually similar static hand poses or dynamic gesture sequences, reducing the dimensionality of high-dimensional hand feature spaces (e.g., raw joint coordinates or image features) for more efficient processing, or detecting anomalous or out-of-distribution gestures. For instance, given a set of unlabeled gesture feature vectors $X=\{x_{1},x_{2},\dots,x_{N}\}$, clustering algorithms, such as K-Means, can partition the data into $K$ distinct groups, by minimizing an objective function like the sum of squared errors within each cluster:
[…]
where $C_{k}$ represents the $k$-th cluster and $\mu_{k}$ its centroid. Such techniques are valuable for initial data exploration, discovering novel gesture categories, or pre-training representations when labeled data is scarce or unavailable.
Self-supervised learning (SSL) is a specialized form of unsupervised learning where supervisory signals are automatically generated from the input data itself by defining "pretext" tasks. In HGR, SSL can learn powerful feature representations from large collections of unlabeled hand images or videos. Common pretext tasks include predicting the temporal order of shuffled video segments of a gesture, reconstructing masked-out patches of a hand image, predicting geometric transformations (e.g., rotation, scaling) applied to a hand image, or employing contrastive learning. In contrastive learning, the model is trained to maximize agreement between different augmented views of the same gesture instance (positive pairs), while minimizing agreement with views from different instances (negative pairs). A widely used objective for this is the InfoNCE loss [204], which for an anchor sample $x_{a}$, its positive counterpart $x_{p}$, and a set of $M$ negative samples $\{x_{n}\}_{n=1}^{M}$, is often formulated as:
$L=-\log\frac{e^{s_{a,p}/\tau}}{e^{s_{a,p}/\tau}+\sum_{n=1}^{M}e^{s_{a,n}/\tau}}$
where $s_{i,j}=\text{sim}(f(x_{i}),f(x_{j}))$, $f(\cdot)$ is an encoder that maps inputs to latent feature vectors, $\text{sim}(\cdot,\cdot)$ denotes a similarity function (e.g., cosine similarity), and $\tau$ is a temperature scaling parameter. The objective encourages the model to assign higher similarity scores to positive pairs relative to negative ones. Representations learned via this mechanism can be effectively transferred and fine-tuned on downstream HGR tasks (e.g., gesture classification or estimation), often yielding notable performance improvements, especially in data-scarce scenarios.
II-F Key Takeaways of this section
This section laid the groundwork by defining fundamental concepts and terminologies essential for understanding Hand Gesture Recognition. The main concepts covered include:
  • Gesture Definition: A gesture is a structured, intentional movement or posture, formally represented as a temporal sequence of poses.
  • Gesture Recognition Tasks: Subdivided into Gesture Classification (mapping input to predefined labels) and Gesture Estimation (digitally representing a gesture’s structure, e.g., pose parameters).
  • Gesture Types: Differentiated into Static Gestures (single, unchanging pose) and Dynamic Gestures (sequence of poses over time).
  • Multimodal Gestures: Involve multiple body parts or sensory modalities (e.g., hand movements with facial expressions or speech) to enhance communication.
  • HGR Focus: This survey specifically targets Hand Gesture Recognition due to its expressiveness, compact interaction zone, precision, and strong research foundation.
  • Learning Paradigms: Beyond supervised learning, Unsupervised Learning (e.g., clustering) and Self-Supervised Learning (SSL) (e.g., contrastive learning with InfoNCE loss) are also used in HGR to leverage unlabeled data.
@@ III-A Article Selection
The selection of articles for this study targets all the recent papers in the computer vision field that specifically focus on hand gesture recognition from images and videos. To ensure a comprehensive review, we began by identifying the top-ranked venues in computer vision as listed under the "Top Publications - Computer Vision & Pattern Recognition" section on Google Scholar [^1] https://scholar.google.com/citations?view_op=top_venues&hl=en&vq=eng_computervisionpatternrecognition. Using a crawler developed for this purpose [^2] https://github.com/manouslinard/gscholar_scraper, we retrieved 2,307 articles from these top venues. Subsequently, we conducted a systematic screening process in which the titles and abstracts of all articles were carefully reviewed based on predefined inclusion criteria detailed in Section III-B, which prioritized recent studies focusing on hand gesture recognition techniques, systems, or applications. This rigorous filtering ensured that only the most relevant to hand gesture recognition contributions from Google Scholar’s top venues for computer vision were included in the survey, and resulted in 37 articles that were directly relevant to the task. Given the limited number of relevant studies (from Google Scholar), we widened the search base by conducting additional searches in all venues listed in Scopus using specific boolean search queries, which are listed in Table II. The queries were carefully designed to contain the most relevant keywords to hand gesture recognition. To ensure consistency with the Google Scholar approach, we systematically reviewed the titles and abstracts of the matching papers from each query, assessing their alignment with the inclusion/exclusion criteria outlined in Section III-B. Articles were excluded if they lacked relevance to the hand gesture recognition task, its techniques, systems, or applications.
TABLE II: Overview of queries addressed to Scopus database and paper counts
@@
"RGB" AND ("hand" OR "gesture") AND ("recognition" OR "classification") 1576 910 24
"sign" AND "gesture" AND "recognition" 2768 1442 32
"Hand" AND ("gesture" OR "pose") AND ("recognition" OR "estimation") AND ("RGB" OR "video" OR "skeleton" OR "multi modal") 2528 746 32
"Multiview" AND "Camera" AND "Hand" AND "Gesture" 5 1 1
Table II includes the following columns:
@@
To determine the values in these columns, we employed a systematic methodology to curate the Scopus papers. The queries were executed sequentially, starting with the first query at the top of the table and progressing to the last query at the bottom. The process began with a broad query focused on terms such as "RGB," "hand," "gesture" and related recognition tasks, providing a foundation for further exploration. This was followed by an expansion into more specific domains, such as sign language recognition, addressed by the second query. To avoid redundancy, articles retrieved from the first query were excluded from the results of the second query. Similarly, the third query broadened the search to include tasks related to hand pose estimation and multimodal approaches, while ensuring duplicates from previous queries were removed. For each collection of retrieved papers, topic modeling using Non-Negative Matrix Factorization (NNMF) was performed in a similar way as detailed in Section III-D. Specifically, NNMF was applied to the titles, keywords, and abstracts of the papers to extract thematic topics. Articles with topics related to the research focus were retained, significantly narrowing down the number of candidate papers. Finally, a selection process was performed based on the inclusion/exclusion criteria outlined in Section III-B, resulting in the final set of papers used in this study.
@@ III-B Inclusion/Exclusion Criteria
  • Studies were included only if published between 2018 and 2025,
@@ III-C Quantitative Analysis
The search and filtering methodology described above helped us identify 125 studies (37 from Scholar and 88 from Scopus) that propose efficient methodologies for hand gesture recognition. To better understand how these studies are distributed over time and across different forums we performed a quantitative analysis, which is summarized in Table III and Fig. 4. The visualizations provide a comprehensive overview of the research landscape of hand gesture recognition, revealing trends in publication frequency, preferred venues, and topics.
@@ III-C1 Distribution of Articles to Venues
As shown in Table III, the main forums where research in hand gesture recognition has been published include the major computer vision conferences such as CVPR, ICCV, ICCVW, ECCV, and the International Journal of Computer Vision. These forums list more than 50% of the relevant works that contribute to advancements in this field.
TABLE III: Distribution of Paper Venues.
@@ III-C2 Research Timeline by Topic
As shown in Fig. 4 the main topics discussed in the articles. The methodology for extracting the topics from the collection of articles is detailed in Section II-D that follows. The bar charts depict the annual occurrences of individual topics, while the line plot represents the total number of publications per year. The analysis reveals a substantial increase in publications in 2024, signaling growing research interest and advancements in the field of hand gesture recognition. The lower number of publications in 2025 is expected, as the year is still ongoing and many papers have yet to be published.
@@ III-D Topic modeling
III-E Key Takeaways of this section
This section detailed the systematic approach undertaken to select, analyze, and categorize the research papers included in this survey. The core aspects of our methodology are:
  • Article Selection: A systematic process was employed, starting with top-ranked computer vision venues (Google Scholar) and expanding with targeted Scopus queries for papers published between 2018-2025.
  • Inclusion/Exclusion Criteria: Focused on studies on hand gesture recognition/estimation from visual input (images/videos), excluding whole-body action recognition without specific hand analysis, and non-English or unavailable publications.
  • Quantitative Analysis: 125 relevant studies were identified. Analysis revealed CVPR, ICCV, ECCV as major publication venues. A significant increase in HGR publications was noted, especially in 2024.
  • Topic Modeling: Non-Negative Matrix Factorization (NNMF) was applied to titles and keywords to identify six primary research themes: Hand Gesture Classification, Hand Gesture Estimation, Sign Language Recognition, Hand/Body Reconstruction, Multimodal Fusion, and Real-Time Recognition.
@@ IV-C1 Skeleton-Based Captures
Skeleton-based methods represent the hand using a skeletal structure that consists of key joints and connections [182, 41, 129]. These captures typically utilize graphs, where nodes represent specific hand joints or segments such as fingers, and edges denote the relationships or connections between them. This graph-based representation enables the system to focus on the underlying structure and movement dynamics of the hand. By abstracting the hand into a simplified skeleton, these methods are particularly effective in scenarios where the precise articulation of fingers and joints is important. An example of an approach using skeleton based capture is shown in Fig. 9.
Figure 9: Example of skeleton-based hand capture. Source [228].
@@ IV-C2 Box/Filter-Based Captures
Box/filter-based methods follow a different approach, focusing on capturing the hand’s region using bounding boxes or regions of interest [7, 31, 157]. These methods involve detecting the hand from the image using a bounding box and once the hand is isolated, applying additional filters or algorithms to extract meaningful features related to the gesture, such as shape, texture, or movement patterns. An example of an approach using box/filter based capture is shown in Fig. 10.
Figure 10: Example of box/filter-based hand capture. Source [180].
The fundamental difference between these approaches lies in how they represent and process hand information. Skeleton-based captures excel in scenarios that require precise motion analysis and are robust against background noise because they directly model the hand’s articulated structure as a set of interconnected joints and segments. This detailed skeletal representation is then typically processed using graph-based algorithms or fed into models designed to understand structural relationships. However, this level of detail often requires more computational resources, both for initially extracting the skeleton and for its subsequent graph-based representation and analysis, making these methods more suitable for applications where accuracy and nuanced articulation are prioritized over raw speed or implementation simplicity.
In contrast, box/filter-based methods adopt a two-stage approach: first, they focus on localizing the hand as a whole within a broader image, typically using a bounding box or region of interest. Once this region is identified, various features (e.g., shape descriptors, texture patterns) are extracted from the pixel data within that box. This makes them well-suited for working with 2D image data and readily adaptable across diverse environments. These methods are generally more user-friendly to implement and are widely applied in hand gesture classification tasks (as described in Section IV-F, where Bayes’ Theorem demonstrates that box/filter methods are predominantly used in classification tasks). Here, the primary goal is to categorize the overall gesture’s appearance or motion pattern, rather than to perform detailed estimation tasks which necessitate precise joint locations and orientations. Despite their advantages in simplicity and directness, box/filter-based methods can be more sensitive to external factors such as inconsistent lighting conditions, cluttered backgrounds that interfere with hand segmentation, and variations in hand appearance, all of which can negatively impact recognition performance by corrupting the features extracted from the hand region.